Need help doing problem 1 and 2
Problem #1: Determine the steady state solution for the given initial conditions: f(t)+2f(t) +5f(t)=0.5; for f(0)=f(t)=0 P
Answer #1

Problem #1: are fcut aft) + 5f(t) = 0.5 $(o)- £10)=0 from above values initial Conditions Lero. does not exist Thus Zero a/2 F(S) - 3+3 6735+23 tractions approach i Paonal S +35+2= - 31 J 9-4ca) - -31 = -1, -2 FCs) = st3 (SH)(sta) 5+3 = A ta 5+1 5+